# Stability of $L^{p}$ Dirichlet solvability under small bi-Lipschitz transformations of domains

### v1

@@ 3.1. Set-up estimates on Lipschitz graph domains
$|K_{12}|\leq\inf_{a\in\mathbb{R}^{n},B\in\mathbb{R}^{(n-1)\times n}}\left|\int_{\mathbb{R}^{n-1}}\nabla_{y}^{2}\eta_{t}\big(y-z\big)\big[h(z)-(a+zB)\big]\,dz\right|\\ \lesssim\inf_{a\in\mathbb{R}^{n},B\in\mathbb{R}^{(n-1)\times n}}\frac{1}{t^{2}}\fint_{B(y,t)}|h(z)-(a+zB)|\,dz=\frac{\beta(y,t)}{t}.$
Then we can argue as in Step 2 starting from (3.28) to obtain that $t|K_{11}|+t|K_{21}|\in\operatorname{CM_{sup}}(C_{M}\epsilon)$. The theorem follows. $\square$
@@ 3.3. Construction of the change of variables
Let $\Omega_{0}:=\{(x,t)\in\mathbb{R}^{n},\,t>g(x)\}$, where $g$ is a $M$-Lipschitz function. There exists $\epsilon_{2}=\epsilon_{2}(M)\in(0,1]$ such that for any $\epsilon\in[0,\epsilon_{2}]$ and for any bi-Lipschitz change of variables $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfying $\|\nabla\Phi-I\|_{\infty}\leq\epsilon$, the map $\rho$ defined in (3.40) is a diffeomorphism between $\Omega_{0}$ and $\Omega:=\Phi(\Omega_{0})$. Moreover, the Jacobian matrix $J$ defined in (3.42) satisfies
@@
(3.44) $|\lambda(X)-h(\pi(X))|\leq Ct(X)\epsilon,$
since $h$ is $\epsilon$-Lipschitz. However, $|y-\pi(X)|=|\pi(X-tv_{n})-\pi(X)|\leq t$, so $|\lambda(X)-h(\pi(X))|\lesssim\epsilon t$ as desired.
We have then, by (3.3), that
(3.45) $|\rho(X)-X-h(\pi(X))|\leq|\lambda(X)-h(\pi(X)|+at|\mathcal{O}-I_{n\times n}|+t|a-1|\lesssim\epsilon t$
by (3.44), $(i)$ of Lemma 3.34, and Lemma 3.37. We now look at the point $X+h(\pi(X))$. Let $z=\pi(X)$ and let $s$ be the distance between $X$ and its projection on $\partial\Omega_{0}$ in the direction of $e_{n}$. Then we can write $X=(z,g(z))+(0,s)$, which shows that
$X+h(\pi(X))=\Phi(z,g(z))+(0,s).$
So if $y\in\mathbb{R}^{n}$, we have
$|X+h(\pi(X))-\Phi(y,g(y))|=|(z-y,g(z)-g(y)+s)+h(z)-h(y)|\\ \geq\sqrt{|z-y|^{2}+|s+g(z)-g(y)|^{2}}-C\epsilon_{2}|z-y|,$
since $h$ is $C\epsilon_{2}$-Lipschitz. If moreover, $|y-z|\geq s/2M$, then of course
$|X+h(\pi(X))-\Phi(y,g(y))|\geq(1-C\epsilon_{2})|z-y|\gtrsim s,$
if $\epsilon_{2}$ is small enough. But if $|y-z|\leq s/2M$, then $|g(z)-g(y)|\leq s/2$, meaning that
$|X+h(\pi(X))-\Phi(y,g(y))|\geq|s|-|g(z)-g(y)|\gtrsim s.$
Altogether, we have $\operatorname{dist}(X+h(\pi(X)),\partial\Omega)\gtrsim s$. With the obvious or previously known estimates, we obtain that
(3.46) $Z:=X+h(\pi(X))\in\mathbb{R}^{n}\setminus\partial\Omega\quad\text{ and }\quad\operatorname{dist}(Z,\partial\Omega)\approx s\approx\operatorname{dist}(X,\partial\Omega_{0})\approx t(X),$
with the implicit constants independent of $\epsilon_{2}\leq 1$. The combination of (3.45) and (3.46) gives that
(3.47) $\rho(X)\in B(Z,C\epsilon_{2}t)\subset B(\rho(X),\operatorname{dist}(X,\partial\Omega)/2)\subset\Omega,$
if $\epsilon_{2}$ is small enough.
Step 2: Decomposition of $J$. Direct computation shows that
@@
as desired.
Step 3: $\rho$ is a diffeomorphism. Note that $\rho$ is smooth on $\Omega_{0}$, because it is essentially constructed from the Green function $G$ which is itself smooth on $\Omega_{0}$. So we just need to prove that $\rho$ is a bijection.
The first key element is that the Jacobian matrix of $\rho$ is invertible and close to identity. Indeed, by (3.42) and (3.33), the Jacobian matrix $\nabla\rho$ of $\rho$ is
@@
where $P$ is the orthogonal matrix that changes the Cartesian base $\{e_{1},\dots,e_{n}\}$ into $\{v_{1},\dots,v_{n}\}$. Step 2 and Lemma 3.34 give us that then
@@
It means that the map $\rho$ is locally invertible at any point. This implies the degree of $\rho$ defined as
$N_{\rho}(\mathcal{Z}):=\#\{X\in\Omega_{0},\,\rho(X)=\mathcal{Z}\}$
is locally constant (positive, integer, not necessarily finite). Therefore $N_{\rho}$ is constant on connected components of $\mathbb{R}^{n}\setminus\partial\Omega$.
We want to show that $N_{\rho}\equiv 1$ on $\Omega$ and $N_{\rho}\equiv 0$ on $\mathbb{R}^{n}\setminus\overline{\Omega}$. One strategy is to construct the family of bi-Lipschitz changes of variables
$\Phi_{\theta}:=\theta\Phi+(1-\theta)I.$
One can see from our explicit construction that $\rho=\rho_{\Phi}$ is continuous in $\Phi$, so $\rho_{\Phi_{\theta}}$ is continuous in $\theta$, and eventually that $N_{\rho_{\theta}}(\Phi_{\theta}(X))\in\mathbb{Z}$ is continuous in $\theta$. Since $\Phi_{0}$ and then $\rho_{0}$ are the identity maps, we have that
$N_{\rho_{\theta}}(\Phi_{\theta}(X))=N_{\rho_{0}}(\Phi_{0}(X))=\left\{\begin{array}[]{ll}1&\text{ if }X\in\Omega_{0}\\ 0&\text{ if }X\in\mathbb{R}^{n}\setminus\overline{\Omega_{0}}.\end{array}\right$.
The theorem follows. $\square$
@@ 4. Proof of the main theorem and corollary
Proof: Take $\epsilon_{2}$ small enough such that Theorem 3.43 applies. Let $P=P(X)$ be the orthogonal matrix that changes the base $\{e_{1},\dots,e_{n}\}$ into $\{v_{1},\dots,v_{n}\}$. Recall that $\nabla\rho$ is given by
@@
By definition, the domain $\Omega_{0,x_{0}}:=\left\{(x,t)\in\mathbb{R}^{n}:\,t>g_{x_{0}}(x)\right\}$ is a convex Lipschitz graph domain, and so Theorem B.6 applies. It means that the $L^{p}$ Dirichlet problem for $\Delta$ is solvable in $\Omega_{0,x_{0}}$, and that the constant $C_{p}$ in (2.13) is independent of $x_{0}$ even though $\Omega_{0,x_{0}}$ itself depends on $x_{0}$. Therefore, by Theorem 1.3, the $L^{p}$ Dirichlet problem for $\Delta$ is solvable in $\Omega_{x_{0}}$. Then Theorem 2.12 entails that there exists a constant $C^{\prime}_{p}:=C^{\prime}_{p}(n,p,C_{p},M)$—so independent of $x_{0}$—such that we have the reverse Hölder bound
@@
Finally note that $\Phi(X)=X+O(\epsilon_{0})|X|$ is close to the identity transform and that $L^{-1}$ is close to $\Psi^{-1}$ in a neighborhood of 0 since $\Psi(X)=LX+o(|X|)$ (using the fact that $\Psi$ is $C^{1}$). Thus by a messy but straightforward computation, we can show that $|\nabla(\Psi\circ\Phi\circ L^{-1})-I|\leq C\epsilon_{0}$ if $r_{0}$ is small enough. This means that $\Psi(\Omega)$ is strongly quasiconvex. ∎
Appendix B Solvability of the $L^{p}$ Dirichlet problem under $C^{1,\alpha}$ diffeomorphism of convex domains

### v2

@@ 3.1. Set-up estimates on Lipschitz graph domains
$|K_{12}|\leq\inf_{a\in\mathbb{R}^{n},B\in\mathbb{R}^{(n-1)\times n}}\left|\int_{\mathbb{R}^{n-1}}\nabla_{y}^{2}\eta_{t}\big(y-z\big)\big[h(z)-a-(z-x)B)\big]\,dz\right|\\ \lesssim\inf_{a\in\mathbb{R}^{n},B\in\mathbb{R}^{(n-1)\times n}}\frac{1}{t^{2}}\fint_{B(y,t)}|h(z)-a-(z-x)B|\,dz=\frac{\beta_{h}(y,t)}{t}.$
We invoke Lemma 3.14 and (3.22), as in Step 2, to obtain that $t|K_{11}|+t|K_{21}|\in\operatorname{CM_{sup}}(C_{M}\epsilon)$. The theorem follows. $\square$
@@ 3.3. Construction of the change of variables
Let $\Omega_{0}:=\{(x,t)\in\mathbb{R}^{n},\,t>g(x)\}$, where $g$ is a $M$-Lipschitz function. There exists $\epsilon_{2}=\epsilon_{2}(M)\in(0,1]$ such that for any $\epsilon\in[0,\epsilon_{2}]$ and for any bi-Lipschitz change of variables $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfying $\|\nabla\Phi-I\|_{\infty}\leq\epsilon$, the Jacobian matrix $J$ defined in (3.41) for the map $\rho$ defined in (3.39) satisfies
@@
Proof: Direct computation shows that
@@
as desired. $\square$
Theorem 3.45.
Let $\Omega_{0}:=\{(x,t)\in\mathbb{R}^{n},\,t>g(x)\}$, where $g$ is a $M$-Lipschitz function. There exists $\epsilon_{2}=\epsilon_{2}(M)\in(0,1]$ such that for any $\epsilon\in[0,\epsilon_{2}]$ and for any bi-Lipschitz change of variables $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfying $\|\nabla\Phi-I\|_{\infty}\leq\epsilon$, the map $\rho$ defined in (3.39) is a diffeomorphism between $\Omega_{0}$ and $\Omega:=\Phi(\Omega_{0})$.
(3.46) $|\lambda(X)-h(\pi(X))|\leq Ct(X)\epsilon,$
since $h$ is $\epsilon$-Lipschitz. However, $|y-\pi(X)|=|\pi(X-tv_{n})-\pi(X)|\leq t(X)$, so we have $|\lambda(X)-h(\pi(X))|\lesssim\epsilon t(X)$ as desired.
We have then, by (3.3), that
(3.47) $|\rho(X)-X-h(\pi(X))|\leq|\lambda(X)-h(\pi(X)|+at|\mathcal{O}-I_{n\times n}|+t|a-1|\lesssim\epsilon t(X)$
by (3.46), $(i)$ of Lemma 3.33, and Lemma 3.36.
We now look at the point $X+h(\pi(X))$. By the definition of $h$, we write
(3.48) $X+h(\pi(X))=X+\Phi\left(\pi(X),g(\pi(X))\right)-\left(\pi(X),g(\pi(X))\right)\\ =\Phi(X)+\Phi\left(\pi(X),g(\pi(X))\right)-\Phi(X)-\left(\pi(X),g(\pi(X))\right)+X.$
We have
$\left|\Phi\left(\pi(X),g(\pi(X))\right)-\Phi(X)-\left(\pi(X),g(\pi(X))\right)+X\right|\leq|X-\left(\pi(X),g(\pi(X))\right)|\left\|\nabla\Phi-I\right\|_{\infty}\\ \leq\epsilon M\operatorname{dist}(X,\partial\Omega_{0})$
since $g$ is $M$-Lipschitz. This, together with (3.47), (3.48) and the estimate $t(X)\approx\operatorname{dist}(X,\partial\Omega_{0})$, gives that
$|\rho(X)-\Phi(X)|\lesssim\epsilon\operatorname{dist}(X,\partial\Omega_{0})\approx\epsilon\operatorname{dist}(\Phi(X),\partial\Omega),$
where in the last inequality we have used that the map $\Phi:\Omega_{0}\to\Omega$ is bi-Lipschitz. Therefore, by choosing $\epsilon_{2}$ sufficiently small, we have that
(3.49) $|\rho(X)-\Phi(X)|\leq\frac{1}{2}\operatorname{dist}(\Phi(X),\partial\Omega),$
which implies that $\rho(X)\in\Omega$. Step 1 follows. Note that (3.49) also implies that
(3.50) $\operatorname{dist}(\rho(X),\partial\Omega)\approx\operatorname{dist}(\Phi(X),\partial\Omega)\operatorname{dist}(X,\partial\Omega_{0}),$
where the second equivalence is due to the fact that $\Phi$ is bi-Lipschitz.
Step 2: We show that $\rho$ is a diffeomorphism. Since $\rho$ is constructed from the Green function $G$, which is smooth on $\Omega_{0}$, the map $\rho$ is smooth on $\Omega_{0}$. Thus, it remains only to prove that $\rho$ is a bijection from $\Omega_{0}$ onto $\Omega$. We shall prove this in 3 steps.
(A) Local invertibility. By (3.41) and (3.32),
@@
where $P$ is the orthogonal matrix that changes the Cartesian base $\{e_{1},\dots,e_{n}\}$ into $\{v_{1},\dots,v_{n}\}$. Theorem 3.42 and Lemma 3.33 give that
@@
It means that the map $\rho$ is locally invertible at any point. More precisely, for every $X\in\Omega_{0}$, there exist $r_{X},r^{\prime}_{X}>0$ such that
(3.53) $\rho_{\theta}|_{B(X,r_{X})}:B(X,r_{X})\to\rho(B(X,r_{X}))\supset B(\rho(X),r^{\prime}_{X})$
is a bijection.
(B) Injectivity and Properness. For $X=(x,t)\in\Omega_{0}$, define the cones
$\operatorname{Cone}(X):=\{(z,s):|z-x|\leq M(s-t)\}.$
Since $g$ is $M$-Lipschitz, for $X,X^{\prime}\in\Omega_{0}$,
$\operatorname{Cone}(X),\operatorname{Cone}(X^{\prime})\subset\Omega_{0}.$
For $X,X^{\prime}\in\Omega_{0}$, choose
$Y_{X}\in\operatorname{Cone}(X)\cap\operatorname{Cone}(X^{\prime})$
with minimal height. The segments $X_{0},Y_{X}]$ and $Y_{X},X]$ lie in $\Omega_{0}$, so by (3.52),
$|\rho(X)-X-\rho(X^{\prime})+X^{\prime}|\leq\big(|X-Y_{X}|+|Y_{X}-X^{\prime}|\big)\|\nabla\rho-I\|_{L^{\infty}(\Omega)}\\ \qquad\leq C_{M}\varepsilon_{2}|X-X^{\prime}|\leq\frac{1}{2}|X-X^{\prime}|,$
provided $\varepsilon_{2}$ is small enough. Consequently,
(3.54) $|\rho(X)-\rho(X^{\prime})|\geq\frac{1}{2}|X-X^{\prime}|\qquad\text{for }X,X^{\prime}\in\Omega_{0},$
which implies that $\rho$ is injective on $\Omega_{0}$.
We claim that for every compact set $K\subset\Omega$,
(3.55) $\rho^{-1}(K)\text{ is compact in }\Omega_{0}.$
Indeed, by (3.50), preimages of compact subsets of $\Omega$ stay away from $\partial\Omega_{0}$. Moreover, since $\rho$ is smooth and so in particular continuous, $\rho^{-1}(K)$ is closed. It therefore suffices to prove that $\rho^{-1}(K)$ is bounded. By fixing $X^{\prime}$ in (3.54), one sees that
(3.56) $|\rho(X)|\to\infty\quad\text{as }|X|\to\infty.$
It means that $\rho^{-1}(K)$ cannot be unbounded, as if it were unbounded, then there would exist a sequence of points $(X_{n})_{n\in\mathbb{N}}\subset\rho^{-1}(K)$ such that $|X_{n}|\to\infty$ as $n\to\infty$. But (3.56) asserts that $|\rho(X_{n})|\to\infty$, contradicting to the assumption that $K$ is compact. So the claim (3.55) follows.
(C) Surjectivity. We now show that $\rho:\Omega_{0}\to\Omega$ is surjective, and thus a diffeomorphism. For $\mathcal{Z}\in\Omega_{0}$, define the degree
$N_{\rho}(\mathcal{Z}):=\#\{X\in\Omega_{0}:\rho(X)=\mathcal{Z}\}.$
We shall prove that
$N_{\rho}(\mathcal{Z})=1\qquad\text{for every }\mathcal{Z}\in\Omega,$
which implies that $\rho$ is a diffeomorphism.
We know from (B) that $\rho$ is injective, and thus $N_{\rho}(\rho(X))=1$. So, we can write
$\Omega:=\underbrace{\{\mathcal{Z}\in\Omega,\,N_{\rho}(\mathcal{Z})=1\}}_{S_{1}}\cup\{\mathcal{Z}\in\Omega,\,N_{\rho}(\mathcal{Z})=0\}.$
In order to prove that $S_{1}=\Omega$, we then just have to show that $S_{1}$ is open and closed in $\Omega$.
$S_{1}$ is open. If $\mathcal{Z}\in S_{1}$, then there exists $X\in\Omega_{0}$ such that $\rho(X)=\mathcal{Z}$, and (3.53) shows that a neighborhood of $\mathcal{Z}$ is contained in the range by $\rho$ of a neighborhood of $X$. Hence $S_{1}$ is open.
$S_{1}$ is closed. Let $(\mathcal{Z}_{k})_{k\in\mathbb{N}}$ be a sequence in $S_{1}$ such that $\mathcal{Z}_{k}\to\mathcal{Z}$ for some $\mathcal{Z}$ as $k\to\infty$. Then we get a sequence $\{X_{k}\}_{k\in\mathbb{N}}$ in $\Omega_{0}$ such that $\rho(X_{k})\to\mathcal{Z}$ as $k\to\infty$. By (3.55), we know that $\{X_{k}\}_{k\in\mathbb{N}}$ is contained in a compact subset of $\Omega_{0}$, hence up to taking a subsequence, we can assume that $X_{k}$ converges to some $X_{\infty}\in\Omega_{0}$. The continuity of $\rho$ gives that $\rho(X_{\infty})=\mathcal{Z}$. We conclude that $\mathcal{Z}\in S_{1}$, which implies that $S_{1}$ is closed.
To summarize, we showed that $\Omega=S_{1}=\{\mathcal{Z}\in\Omega,\,N_{\rho}(\mathcal{Z})=1\}$, which means that the local diffeomorphism $\rho$ is a (global) diffeomorphism from $\Omega_{0}$ to $\Omega$. The theorem follows. $\square$
@@ 4. Proof of the main theorem and corollary
Proof: Take $\epsilon_{2}$ small enough such that Theorems 3.42 and 3.45 apply. Let $P=P(X)$ be the orthogonal matrix that changes the base $\{e_{1},\dots,e_{n}\}$ into $\{v_{1},\dots,v_{n}\}$. Recall that $\nabla\rho$ is given by
@@
By definition, the domain $\Omega_{0,x_{0}}:=\left\{(x,t)\in\mathbb{R}^{n}:\,t>g_{x_{0}}(x)\right\}$ is a convex Lipschitz graph domain, and so Theorem C.6 applies. It means that the $L^{p}$ Dirichlet problem for $\Delta$ is solvable in $\Omega_{0,x_{0}}$, and that the constant $C_{p}$ in (2.13) is independent of $x_{0}$ even though $\Omega_{0,x_{0}}$ itself depends on $x_{0}$. Therefore, by Theorem 1.3, the $L^{p}$ Dirichlet problem for $\Delta$ is solvable in $\Omega_{x_{0}}$. Then Theorem 2.12 entails that there exists a constant $C^{\prime}_{p}:=C^{\prime}_{p}(n,p,C_{p},M)$—so independent of $x_{0}$—such that we have the reverse Hölder bound
@@
Finally note that $\Phi(X)=X+O(\epsilon_{0})|X|$ is close to the identity transform and that $L^{-1}$ is close to $\Psi^{-1}$ in a neighborhood of 0 since $\Psi(X)=LX+o(|X|)$ (using the fact that $\Psi$ is $C^{1}$). Thus by the chain rule, we can show that $|\nabla(\Psi\circ\Phi\circ L^{-1})-I|\leq C\epsilon_{0}$ if $r_{0}$ is small enough. This means that $\Psi(\Omega)$ is strongly quasiconvex. ∎
Appendix B An alternative proof of Lemma 3.2
We can assume without loss of generality that $g(0)=0$, that is, $0\in\partial\Omega_{0}$.
$G^{Y}(X):=\int_{\Omega_{0}}\eta_{\delta(Y)}(Y-Z)G(X,Z)\,dZ.$
[…]
where $K^{*}:=\bigcup_{Z\in K}\overline{B(Z,\delta(Z)/2)}$, and the final convergence is by (B.2).
We claim that
[…]
Integration by parts in $Z_{n}$ gives that
Consequently, $\partial_{n}G^{Y}+H^{Y}-u_{\partial_{n}G^{Y}}$ is a function with non-tangential bounds and zero trace, and is a solution to $\Delta u=0$ in $\Omega_{0}$ by (B.4). Uniqueness of the Dirichlet problem (see [Ken94, Theorem 1.4.4], or [HLMP22, Section 6] for unbounded domains) gives the formula:
$\partial_{n}G^{Y}=u_{\partial_{n}G^{Y}}-H^{Y}\quad\text{in }\Omega_{0}.$
$\partial_{n}G^{Y}(x)=\lim_{h\to 0\atop h>0}\frac{G^{Y}(x+he_{n})}{h}\geq 0.$
Second, we compare $H^{Y}$ and $G^{Y}$ outside of $B_{Y}$. Observe that
by integration by parts. Therefore,
$|H^{Y}(X)|\leq\int_{B_{Y}}\eta_{\delta(Y)}(Y-Z)|\nabla_{Z}G(X,Z)|dZ.$
If $B_{X}$ and $B_{Y}$ do not intersect, we apply Lemma 2.19 to $G(X,\cdot)$ in $\Omega_{0}\setminus B_{X}$. For $Z\in B_{Y}$, $\operatorname{dist}(Z,\partial(\Omega_{0}\setminus B_{X}))=\min\left\{\delta(Z),\operatorname{dist}(Z,B_{X})\right\}$. Obviously, $\delta(Z)\geq\delta(Y)/2$. We check that we have $\operatorname{dist}(Z,B_{X})\geq\max\left\{\delta(X)/2,|Y-X|-\delta(X)/2-\delta(Y)/2\right\}$. If $\delta(X)<\delta(Y)/4$, then $|Y-X|\geq\delta(Y)-\delta(X)$, and so $|Y-X|-\delta(X)/2-\delta(Y)/2\geq\delta(Y)/8$. This implies that $\operatorname{dist}(Z,B_{X})\geq\delta(Y)/8$, and thus for $Z\in B_{Y}$, $\operatorname{dist}(Z,\partial(\Omega_{0}\setminus B_{X}))\geq\delta(Y)/8$. So we have that
(B.5) $|H^{Y_{i}}(X)|\lesssim 2^{-i}G^{Y_{i}}(X)\lesssim cG^{Y_{i}}.$
For $c>0$, using successively Step 1a, Step 1b, the positivity of $u_{\partial_{n}G^{Y_{i}}}$ and (B.5) we have then that
Since the bound is for all $c>0$, we conclude that
$\frac{\partial_{n}G(X)}{G(X)}\geq 0$
as desired.
(B.6) $\partial_{n}G>0.$
(B.7) $\partial_{n}G(X)\geq cG(X)/\delta(X)\qquad\text{ for all }X\in\Omega_{0}.$
This claim easily yields (H1)–(H2).
To show (B.7), it suffices to show that there is $c>0$ depending only on $n$ and the Lipschitz constant of $g$, such that $\partial_{n}G(0,1)>cG(0,1)$ (recall that $0\in\partial\Omega_{0}$). We can assume without loss of generality that $G(0,1)=1$. By the Harnack inequality, $0<G(X)\lesssim 1$ for all $X\in B(0,2)\cap\Omega_{0}$, with the implicit constant depending only on the dimension and the Lipschitz constant. Moreover, since $G$ is smooth (and in particular, continuous) in $B(0,1)\cap\Omega_{0}$, there exists $c_{0}>0$ (depending only on $n$ and the Lipschitz constant of $g$) such that $G(0,c_{0})\leq 1/2$. So we have that
$\frac{1}{2}\leq G(0,1)-G(0,c_{0})=\int_{c_{0}}^{1}\partial_{n}G(0,s)ds.$
By the mean value theorem, there exists $s_{0}\in(c_{0},1)$ such that $\partial_{n}G(0,s_{0})\geq 1/2$. Finally, since $\partial_{n}G$ is harmonic and nonnegative, the Harnack inequality implies that $\partial_{n}G(0,1)\geq c$ for some $c>0$.
Appendix C Solvability of the $L^{p}$ Dirichlet problem under $C^{1,\alpha}$ diffeomorphism of convex domains